# Analyzing Message-Code Inconsistency in AI Coding Agent-Authored Pull Requests

### v1

@@ Abstract.
$c$s: Software and its engineering Collaboration in software development
@@ 1. Introduction
  1. (1)
    Annotated dataset: We release 974 manually annotated PRs (432 partial/misaligned) drawn from AIDev to support the development of more reliable AI coding agents.
  2. (2)
    Prevalence (RQ1): We found 406 high-MCI PRs (1.7%) with a 20-fold variation across agents, suggesting that a non-trivial fraction of PRs could mislead reviewers.
  3. (3)
    Taxonomy (RQ2): We identified eight PR-MCI types, finding that Phantom Changes (descriptions claim unimplemented changes) dominated (45.4%).
  4. (4)
@@ 2. Background and Related Work
Message-Code Inconsistency. Commit messages and PR descriptions serve as documentation of developer intent. When these descriptions misalign with actual code changes, they create message-code inconsistency (MCI) (Wang et al., 2025a; Dong et al., 2023). Recent work has studied MCI in human-written commit messages and benchmarked LLMs for detecting text-code inconsistencies (Dong et al., 2023; Zhang et al., 2025), highlighting challenges in generating faithful change descriptions.
Gap in PR-MCI. While MCI has been studied for human-written commit messages (Dong et al., 2023; Zhang et al., 2025), and prior work has analyzed PR text for quality concerns (Karmakar et al., 2022), very few prior work examines inconsistency in Agentic-PR descriptions. This gap is critical because agents generate descriptions automatically without human oversight, potentially producing misleading explanations at scale. PR decision-making research shows that description quality is associated with acceptance rates (Zhang et al., 2022), and automatically created PRs can differ in how maintainers interact with them (Wyrich et al., 2021). Our work extends this research by examining how description-code consistency is associated with outcomes for Agentic-PRs. Note that PR descriptions differ from commit messages: PR descriptions may span multiple commits and serve as the primary communication channel for reviewers.
@@ 3. Methodology and Experiment Setup
We structured the study around three research questions (RQs):
  • RQ1 (Prevalence): How frequently do Agentic-PR descriptions misalign with their code changes?
  • RQ2 (Taxonomy): What types of message-code inconsistencies occur in AI-generated PRs?
  • RQ3 (Impacts): Is message-code inconsistency associated with PR acceptance and review effort?
Figure 1 illustrates our workflow. We used the AIDev dataset (Li et al., 2025) and analyzed the AIDev-pop subset (33,596 PRs from 2,807 repositories with >>100 stars) via the pull_request split. After filtering for closed PRs, permissive licenses (MIT or Apache 2.0), and the six most common task types (95.2%), the final dataset contains 23,247 PRs authored by five AI agents.
@@ 3.1. Measuring PR-MCI (RQ1)
Following prior MCI research (Wen et al., 2019; Zhang et al., 2025), we used PR message-code inconsistency (PR-MCI) as the degree to which a PR description (title + body) diverges from the underlying code changes. We measured PR-MCI using a heuristic similarity score that combines three complementary signals: scope adequacy $s_{s}$ (whether description verbosity matches code churn), file-type consistency $s_{f}$ (whether mentioned file types, e.g., tests or documentation, are actually modified), and task-type alignment $s_{t}$ (whether description language matches the labeled task type). The similarity score $s\in[0,1]$ aggregates these signals using a weighted scheme (Kaliszewski and Podkopaev, 2016):
(1) $s=0.3\cdot s_{s}+0.4\cdot s_{f}+0.3\cdot s_{t}$
where $s_{f}$ receives the highest weight (40%) for providing the most concrete and verifiable evidence of misalignment (e.g., claiming test changes when no test files are modified), while $s_{s}$ and $s_{t}$ each receive 30% as complementary but less definitive signals (Barnett et al., 2016).
To validate the PR-MCI score and calibrate the decision threshold, two authors manually annotated 600 PRs, stratified by agent, task type, and score bins, as aligned, partially aligned, or misaligned, achieving strong agreement ($\kappa=0.892$). PRs with similarity below $\theta=0.61$ are labeled high-MCI, where $\theta$ is selected by optimizing F1 on labeled validation data (Fan and Lin, 2007) after conducting sensitivity analysis across threshold ranges (Lipton et al., 2014); five-fold cross-validation shows stable threshold selection (mean=0.606, std=0.008) (Zhang et al., 2025). We also evaluated alternative methods: an embedding model (Qwen3-Embedding-0.6B (Qwen Team, 2024)) performs poorly (F1=0.150), and an agreement method (both methods must agree) achieves lower F1 (0.567 vs. 0.630), indicating both are insufficient (see Table 1).
Consistent with large-scale inconsistency studies (Wen et al., 2019), we computed PR-MCI automatically for all PRs and reported prevalence with 95% Wilson confidence intervals.
@@ 3.3. Statistical Tests (RQ3)
To examine associations between PR-MCI and outcomes, we compared high- and low-MCI PRs using standard statistical tests. Following PR decision research (Zhang et al., 2022), we analyzed acceptance rate, review count, comment count, and time to merge. We applied Chi-square tests for binary outcomes and Mann-Whitney U tests for skewed continuous variables, reported Cramér’s V and Cliff’s $\delta$ as effect sizes. Significance is assessed at $p<0.05$.
@@ 4.1. RQ1: Prevalence and Patterns of PR-MCI
Table 1 reports validation metrics and prevalence by agent and task type. Overall, 406 out of 23,247 Agentic-PRs (1.7%) exhibited high PR-MCI (similarity score below 0.61), where descriptions poorly match code changes, suggesting that a small but non-trivial fraction could mislead reviewers. The prevalence revealed a 20-fold difference across agents: GitHub Copilot had the highest rate (8.7%, 234 high-MCI PRs out of 2,675), followed by Cursor (4.5%, 31 out of 682). Sensitivity analysis (threshold 0.55-0.65) suggested that relative ordering remained consistent across thresholds  [^1] Sensitivity analysis results are available in our replication package..
@@ 4.2. RQ2: PR-MCI Taxonomy
Figure 3 shows the overall taxonomy, by agent, and by task type. Overall, the most common PR-MCI type was Phantom Changes (45.4%), where PR descriptions claim unimplemented changes. This was followed by Scope Understated (22.0%), where descriptions omit significant changes, and Placeholder/Incomplete (18.8%), which used generic or boilerplate text [^2] Examples of different PR-MCI types are provided in our replication package..
For agent-specific patterns, GitHub Copilot was dominated by Phantom Changes (74.0%), while Cursor (52.8%) and Devin (41.0%) showed Scope Understated as the primary issue. OpenAI Codex showed a balanced distribution with Placeholder/Incomplete (39.8%) and Scope Understated (41.9%). Task-specific patterns showed that Test (55.0%) and Bug Fix (53.5%) PRs were most prone to Phantom Changes, while Refactor PRs showed Scope Understated (40.5%) as the dominant issue.
@@
4.3. RQ3: Association of PR-MCI with Review and Outcome
@@
Metric Low-MCI High-MCI $p$-value Effect Size # (Low) # (High) Acceptance Rate (%) 80.0 28.3 <<0.001 Small 22,841 406 Mean Time to Merge (h) 16.0 55.8 <<0.001 Small 18,282 115 Mean # Review 0.74 0.42 0.039 Negligible 22,841 406 Mean # Comment 0.64 0.37 0.012 Negligible 22,841 406 Acceptance Rate by Agent (%) Claude Code 69.3 50.0 1.000 Negligible 218 2 GitHub Copilot 59.4 3.4 <<0.001 Medium 2,441 234 Cursor 73.3 67.7 0.638 Negligible 651 31 Devin 58.3 45.5 0.580 Negligible 2,880 11 OpenAI Codex 87.2 62.5 <<0.001 Negligible 16,651 128 Acceptance Rate by Task Type (%) Chore 80.2 45.8 <<0.001 Small 576 24 Documentation 89.8 34.5 <<0.001 Small 2,786 29 Feature 80.1 27.7 <<0.001 Small 10,782 166 Bug Fix 73.6 17.7 <<0.001 Small 5,206 113 Refactor 78.1 42.6 <<0.001 Small 1,499 54 Test 84.3 25.0 <<0.001 Small 1,992 20 Time to Merge (h) by Agent (merged PRs only) Claude Code 31.4 6.1 0.817 Negligible 149 1 GitHub Copilot 70.2 24.0 0.163 Small 1,420 8 Cursor 23.1 28.9 0.333 Negligible 477 21 Devin 21.8 14.4 0.976 Negligible 1,669 5 OpenAI Codex 5.7 38.6 <<0.001 Medium 14,509 77 Time to Merge (h) by Task Type (merged PRs only) Chore 16.5 81.4 0.181 Small 460 11 Documentation 8.9 1.1 0.140 Small 2,493 10 Feature 11.3 16.4 <<0.001 Medium 8,617 44 Bug Fix 20.6 34.4 <<0.001 Medium 3,812 20 Refactor 15.5 51.2 0.012 Small 1,165 22 Test 6.4 80.6 0.825 Negligible 1,677 5
Table 2 reports summary statistics for key metrics with statistical tests and breakdowns by agent and task type. Notably, high-MCI PRs were associated with a significantly lower acceptance rate: 28.3% vs. 80.0%, a difference of 51.7% ($p<0.001$). Effect size was small (Cramér’s V = 0.166), indicating a non-trivial practical difference. The relationship varied substantially by agent: for GitHub Copilot, the effect was largest (3.4% vs. 59.4%, 55.9 percentage point difference, $p<0.001$). The effect of high-MCI was even more obvious by task type, with statistically significant differences for all tasks (all $p<0.001$ with small effect sizes).
For merged PRs, high-MCI PRs took significantly longer to merge: 55.8 hours vs. 16.0 hours ($p<0.001$; Cliff’s $\delta=0.310$). The effect was most pronounced for OpenAI Codex (38.6 vs. 5.7 hours, $p<0.001$) and refactoring tasks (51.2 vs. 15.5 hours, $p=0.012$).
Further, regression analysis [^3] Detailed regression results are available in our replication package. showed that PR-MCI remained significantly associated with lower acceptance and longer merge times ($p<0.001$ for both) after controlling for code churn, files changed, task type, and agent, suggesting the association is not fully explained by code complexity or task characteristics alone.
@@ 5.4. Threats to Validity
This study has several limitations that should be considered when interpreting the results. (1) Construct validity: Our PR-MCI metric is a heuristic proxy for semantic consistency and may miss subtle mismatches or penalize concise but accurate descriptions. (2) Threshold calibration: The decision threshold is calibrated on a validation set, which weakens absolute prevalence claims despite observed stability under cross-validation. (3) External validity: Dataset filtering choices prioritize scale over strict controls and may affect generalizability. (4) Sample size imbalance: Agent sample sizes are imbalanced (e.g., Claude Code has only 220 PRs), limiting the reliability of agent-specific conclusions for smaller samples. (5) Taxonomy completeness: The taxonomy is derived from a limited subset of PRs (432 partial/misaligned cases) and may not capture all inconsistency types. (6) Causal inference: Our analysis is observational and does not establish causal relationships between PR-MCI and review outcomes.
5.5. Ethical Considerations
This study analyzes only publicly available GitHub data from repositories under permissive licenses (MIT or Apache 2.0) and reported findings in aggregate to avoid identifying individuals or projects.
@@ 6. Conclusion
Data Availability. All data, scripts, results, extra analysis, and supplementary materials are available in our replication package.

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: 23rd International Conference on Mining Software Repositories; April 13–14, 2026; Rio de Janeiro, Brazilbooktitle: 23rd International Conference on Mining Software Repositories (MSR ’26), April 13–14, 2026, Rio de Janeiro, Brazildoi: 10.1145/3793302.3793583isbn: 979-8-4007-2474-9/2026/04ccs: Software and its engineering Collaboration in software development
@@ 1. Introduction
(1) Annotated dataset: We release 974 manually annotated PRs (432 partial/misaligned cases), supporting the development of PR-MCI detection methods and more reliable AI coding agents.
(2) RQ1: How often do Agentic-PR descriptions misalign with their code changes? We found 406 high-MCI PRs (1.7%) with a 20-fold variation across agents, suggesting that a non-trivial fraction of PRs could mislead reviewers.
(3) RQ2: What types of message-code inconsistencies occur in AI-generated PRs? We identified eight PR-MCI types, finding that Phantom Changes (descriptions claim unimplemented changes) dominated (45.4%).
@@ 2. Background and Related Work
Message-Code Inconsistency. Commit messages and PR descriptions serve as documentation of developer intent. When these descriptions misalign with actual code changes, they create message-code inconsistency (Wang et al., 2025a; Dong et al., 2023). Recent work has studied MCI in human-written commit messages and benchmarked LLMs for detecting text-code inconsistencies (Dong et al., 2023; Zhang et al., 2025), highlighting challenges in generating faithful change descriptions.
Gap in PR-MCI. While MCI has been studied for human-written commit messages (Dong et al., 2023; Zhang et al., 2025), and prior work has analyzed PR text for quality concerns (Karmakar et al., 2022), very few prior works examine inconsistency in Agentic-PR descriptions [^1] Note that PR descriptions differ from commit messages: PR descriptions may span multiple commits and serve as the primary communication channel for reviewers.. This gap is critical because agents generate descriptions automatically without human oversight, potentially producing misleading explanations at scale. PR decision-making research shows that description quality is associated with acceptance rates (Zhang et al., 2022), and automatically created PRs can differ in how maintainers interact with them (Wyrich et al., 2021). Our work extends this research by examining how description-code consistency is associated with outcomes for Agentic-PRs.
@@ 3. Methodology and Experiment Setup
Figure 1 illustrates our workflow. We used the AIDev-pop subset (33,596 PRs from 2,807 repositories with >>100 stars) from the AIDev dataset (Li et al., 2025). After filtering for closed PRs, permissive licenses (MIT or Apache 2.0), and six most common task types (95.2%), the final dataset contains 23,247 PRs authored by five AI agents.
@@ 3.1. Measuring PR-MCI (RQ1)
Following prior MCI research (Wen et al., 2019; Zhang et al., 2025), we used PR-MCI as the degree to which a PR description (title + body) diverges from the underlying code changes. We measured PR-MCI using a heuristic similarity score that combines three complementary signals: scope adequacy $s_{s}$ (whether description verbosity matches code churn), file-type consistency $s_{f}$ (whether mentioned file types, e.g., tests or documentation, are actually modified), and task-type alignment $s_{t}$ (whether description language matches the labeled task type). The similarity score $s\in[0,1]$ aggregates these signals using a weighted scheme (Kaliszewski and Podkopaev, 2016): $s=0.3\cdot s_{s}+0.4\cdot s_{f}+0.3\cdot s_{t}$, where $s_{f}$ receives the highest weight (40%) for providing the most concrete and verifiable evidence of misalignment (e.g., claiming test changes when no test files are modified), while $s_{s}$ and $s_{t}$ each receive 30% as complementary but less definitive signals (Barnett et al., 2016)  [^2] Component ablation analysis is available in our supplementary material..
To validate the PR-MCI score and calibrate the decision threshold, two authors manually annotated 600 PRs, stratified by agent, task type, and score bins, as aligned, partially aligned, or misaligned, achieving strong agreement ($\kappa=0.892$). PRs with similarity below $\theta=0.61$ are labeled high-MCI, where $\theta$ is selected by optimizing F1 on labeled validation data (Fan and Lin, 2007) after conducting sensitivity analysis across threshold ranges (Lipton et al., 2014); five-fold cross-validation shows stable threshold selection (mean=0.606, std=0.008) (Zhang et al., 2025). We also evaluated alternative methods: an embedding model (Qwen3-Embedding-0.6B (Qwen Team, 2024)) performs poorly (F1=0.150), and an agreement method (both methods must agree) achieves lower F1 (0.567 vs. 0.630), indicating both are insufficient (see Table 1). Consistent with large-scale inconsistency studies (Wen et al., 2019), we report prevalence with 95% Wilson confidence intervals.
@@ 3.3. Statistical Tests (RQ3)
To examine associations between PR-MCI and outcomes, we compared high- and low-MCI PRs using standard statistical tests. Following PR decision research (Zhang et al., 2022), we analyzed acceptance rate, time to merge, review count, and comment count. For binary outcomes (acceptance), we used Chi-square tests (McHugh, 2013) with Cramér’s V (Cramér, 1999); for continuous variables (e.g. time), Mann-Whitney U tests (McKnight and Najab, 2010) with Cliff’s $\delta$ (Macbeth et al., 2011). Significance is assessed at $p<0.05$.
@@ 4.1. RQ1: Prevalence of PR-MCI
Table 1 reports prevalence and validation metrics. Overall, 406 out of 23,247 Agentic-PRs (1.7%) exhibited high PR-MCI (similarity score below 0.61), where descriptions poorly match code changes, suggesting that a small but non-trivial fraction could mislead reviewers. The prevalence revealed a 20-fold difference across agents: GitHub Copilot had the highest rate (8.7%, 234 high-MCI PRs out of 2,675), followed by Cursor (4.5%, 31 out of 682). [^3] Sensitivity analysis (threshold 0.55-0.65) suggested that relative ordering remained consistent across thresholds, as in our supplementary material.
@@ 4.2. RQ2: PR-MCI Taxonomy
Figure 3 shows the taxonomy over all, by agent, and by task type. In particular, the most common PR-MCI type was Phantom Changes (45.4%), where PR descriptions claim unimplemented changes. This was followed by Scope Understated (22.0%), where descriptions omit significant changes, and Placeholder/Incomplete (18.8%), which used generic or boilerplate text [^4] Examples of different PR-MCI types are provided in our supplementary material..
For agent-specific patterns, GitHub Copilot was dominated by Phantom Changes (74.0%), while Scope Understated was the primary issue for Cursor (52.8%) and Devin (41.0%). OpenAI Codex showed a balanced distribution with Placeholder/Incomplete (39.8%) and Scope Understated (41.9%).
Task-specific patterns showed that Test (55%) and Bug Fix (53.5%) PRs were most prone to Phantom Changes, while Refactor PRs showed Scope Understated (40.5%) as the dominant issue.
@@
4.3. RQ3: Impact on Review Outcomes
@@
Metric Low-MCI High-MCI $p$-value Effect Size # Low # High Acceptance Rate (%) 80.0 28.3 <<0.001 Small 22,841 406 Mean Time to Merge (h) 16.0 55.8 <<0.001 Small 18,282 115 Mean # Review 0.74 0.42 0.039 Negligible 22,841 406 Mean # Comment 0.64 0.37 0.012 Negligible 22,841 406 Acceptance Rate by Agent (%) Claude Code 69.3 50.0 1.000 Negligible 218 2 GitHub Copilot 59.4 3.4 <<0.001 Medium 2,441 234 Cursor 73.3 67.7 0.638 Negligible 651 31 Devin 58.3 45.5 0.580 Negligible 2,880 11 OpenAI Codex 87.2 62.5 <<0.001 Negligible 16,651 128 Acceptance Rate by Task Type (%) Chore 80.2 45.8 <<0.001 Small 576 24 Documentation 89.8 34.5 <<0.001 Small 2,786 29 Feature 80.1 27.7 <<0.001 Small 10,782 166 Bug Fix 73.6 17.7 <<0.001 Small 5,206 113 Refactor 78.1 42.6 <<0.001 Small 1,499 54 Test 84.3 25.0 <<0.001 Small 1,992 20 Time to Merge (h) by Agent (merged PRs only) Claude Code 31.4 6.1 0.817 Negligible 149 1 GitHub Copilot 70.2 24.0 0.163 Small 1,420 8 Cursor 23.1 28.9 0.333 Negligible 477 21 Devin 21.8 14.4 0.976 Negligible 1,669 5 OpenAI Codex 5.7 38.6 <<0.001 Medium 14,509 77 Time to Merge (h) by Task Type (merged PRs only) Chore 16.5 81.4 0.181 Small 460 11 Documentation 8.9 1.1 0.140 Small 2,493 10 Feature 11.3 16.4 <<0.001 Medium 8,617 44 Bug Fix 20.6 34.4 <<0.001 Medium 3,812 20 Refactor 15.5 51.2 0.012 Small 1,165 22 Test 6.4 80.6 0.825 Negligible 1,677 5
Table 2 reports summary statistics for key metrics with statistical tests and breakdowns by agent and task type. Notably, high-MCI PRs were associated with a significantly lower acceptance rate: 28.3% vs. 80.0%, a difference of 51.7% ($p<0.001$). The corresponding effect size was small (Cramér’s V = 0.166), indicating a non-trivial practical difference. The impacts on acceptance rate varied substantially by agent: for GitHub Copilot, the effect was largest (3.4% vs. 59.4%, a 55.9 percentage point drop, $p<0.001$). The effect of high-MCI was even more obvious by task type, with statistically significant differences for all tasks (all $p<0.001$ with small effect sizes).
For the time to merge, high-MCI PRs took significantly longer: 55.8 hours vs. 16.0 hours ($p<0.001$; small effect size with Cliff’s $\delta=0.310$). The impact was most pronounced for OpenAI Codex (38.6 vs. 5.7 hours, $p<0.001$) and refactoring tasks (51.2 vs. 15.5 hours, $p=0.012$).
Further, regression analysis [^5] Detailed regression results are available in our supplementary material. showed that PR-MCI remained significantly associated with lower acceptance and longer merge times ($p<0.001$ for both) after controlling for code churn, files changed, task type, and agent, suggesting that PR-MCI has an independent association with outcomes beyond what can be explained by code complexity or task characteristics.
@@ 5.4. Threats to Validity
This study has several limitations that should be considered when interpreting the results. (1) Construct validity: Our PR-MCI metric is a heuristic proxy for semantic consistency and may miss subtle mismatches or penalize concise but accurate descriptions. (2) Threshold calibration: The decision threshold is calibrated on a validation set, which weakens absolute prevalence claims despite observed stability under cross-validation. (3) External validity: Dataset filtering choices prioritize scale over strict controls and may affect generalizability. (4) Sample size imbalance: Agent sample sizes are imbalanced (e.g., Claude Code has only 220 PRs), limiting the reliability of agent-specific conclusions for smaller samples. (5) Taxonomy completeness: The taxonomy is derived from a limited subset of PRs (432 partial/misaligned cases) and may not capture all inconsistency types. (6) Causal inference: Our analysis is observational and does not establish causal relationships between PR-MCI and review outcomes. (7) Ethical considerations: This study analyzes only publicly available GitHub data from repositories under permissive licenses (MIT or Apache 2.0) and reported findings in aggregate to avoid identifying individuals or projects.
@@ 6. Conclusion
Data Availability. All data, scripts, results, and supplementary materials are available at https://github.com/gjz78910/PR-MCI.
Acknowledgment. This work has been supported by the ITEA GENIUS grant (project number 23026).